# A Forward Reachability Perspective on Control Barrier Functions and Discount Factors in Reachability Analysis

### v3

@@ 5 Discussion
Although inspired by existing works that establish the connection between reachability and robust control invariance, our work is the first paper that connects forward reachability to the analysis of robust control invariance. Existing works have focused on backward reachability-based formulations that produce the largest robust control invariant set contained in a given desired safety region called the viability kernel. This is done by finding the inevitable (or minimal) backward reachable tube (BRT) [27] of the unsafe region, which becomes the complement of the viability kernel. Various value functions have been proposed to characterize the viability kernels by computing the BRTs, both for finite [13, 26, 28, 19] and infinite time [17, 38, 2, 35] but none of these functions are CBFs, as indicated in Table 1.
The time-varying value functions that characterize finite-horizon BRTs are proposed in [26, 28, 19], where the discount is not needed for the boundedness and continuity of the value function, and the solution uniqueness of the HJ-PDE. The work in [17] has extended these formulations to the infinite-horizon setting (Table 1, first row). However, the value function can be discontinuous and the corresponding HJ-PDE admits non-unique solutions [20]. The work in [2, 38] proposes formulations that resolve these issues for the infinite-horizon setting through the introduction of the discount factor. However, the differential inequality condition on the value function that emerges from the corresponding HJ-PDE differs from the barrier constraint (Table 1, second row). Moreover, the value function flattens to zero in the interior of the control invariant set, thus, there is no non-zero gradient of the value function inside the control invariant set that can be useful for the synthesis of safety control. Thus, this value function is distinct from CBFs, and the authors in [39] extend this formulation to a new notion called Guidance-barrier functions. On the other hand, the work in [13] presents a formulation for the finite-horizon BRT wherein the differential inequality derived from the corresponding HJ-PDE matches the barrier constraint in the CBF definition. However, its extension to infinite horizon might lead to an unbounded and discontinuous value function and non-unique solutions to the corresponding HJ-PDE (Table 1, third row).
Our formulation of forward reachability ensures compliance with the barrier constraint in the CBF definition. Additionally, the value function is both continuous and bounded in $\mathbb{R}^{n}$, while the corresponding HJ-PDE has a unique solution. The central idea behind our approach is the usage of the discount factor backward in time, as $e^{\gamma t}$ where $t<0$, in the definition of the discounted FRT value function (18). In contrast to the discount in BRT formulations leading to the emergence of $-\gamma V_{\gamma}(x)$ in the corresponding HJ-PDEs (Table 1, second row), the usage of discount in this way leads to the emergence of positive $\gamma V_{\gamma}(x)$ term in the FRT-HJ-VI (21), and thus the satisfaction of the barrier constraint (Table 1, last row). Moreover, $e^{\gamma t}$ vanishes as $t$ approaches $\infty$- ∞, thereby ensuring continuous, bounded value functions and the solution uniqueness of the FRT-HJ-VI, resulting from the contraction mapping property outlined in Section 4.2.
An example in Appendix 7.8 illustrates that previous approaches cannot satisfy the barrier constraint, continuity, and boundedness all at the same time, whereas our formulation does.
@@ 6 Conclusions
As we look toward future research avenues, several open questions and challenges emerge. One salient assumption underlying our study is the differentiability of the boundaries of control invariant sets. A deeper understanding of the implications and limitations of this assumption is crucial for broadening the applicability of our results. For example, the relevance of such regularity properties of the boundary of the safe set for stochastic systems is elucidated in [11]. Moreover, the potential of forward reachability, especially in the context of inevitable FRT, has been discussed but not yet fully explored. For instance, our finding of the contraction mapping of the discounted FRT Bellman backup has potential ramifications for learning-based approaches. This property may pave the way for advancements in value-function-based approximate dynamic programming algorithms for safety control [18, 7].
@@ 7.1 Proof of Lemma 7
$\rho_{\mathrm{u}}(\mathrm{u})(-t)=\mathrm{u}(t),\quad\forall t\in[0,\infty);\vspace{-0.5em}$ (30)
@@
$\rho_{\xi_{d}}(\xi_{d})[\mathrm{u}^{-}](-t)=\xi_{d}[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t),\vspace{-0.5em}$ (31)
@@
$\mathrm{x}^{-}(-t)=\mathrm{x}(t)\quad\forall t\in[0,\infty),\vspace{-0.5em}$ (32)
@@
$\dot{\mathrm{x}}(t)=-f(\mathrm{x}(t),\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})(t),\rho_{\xi_{d}}^{-1}(\Xi_{d}^{-})[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t))\vspace{-0.25em}$
for $t>0$, and $\mathrm{x}(0)=x$. Since $\rho_{\mathrm{u}}$ and $\rho_{\xi_{d}}$ are one-to-one, the viability kernel of $\mathrm{Int}(S)^{c}$ under $-f$, defined in Definition 7, is equivalent to the following set:
[…]
$\forall t\in[0,T],\mathrm{x}(t)\in C+B_{\varepsilon}\},\vspace{-0.5em}$ (33)
@@
By [38, Lemma 1], the viability kernel (33) of $\mathrm{Int}(S)^{c}$ is characterized by a particular value function:
[…]
$\mathrm{FRT}(\mathrm{Int}(S))=\{x~{}|~{}V_{\gamma}(x)\neq 0\}.\vspace{-0.5em}$ (35)
Since $h_{S}$ is bounded,
$V_{\gamma}(x)\geq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}[e^{\gamma t}h_{S}(\mathrm{x}^{-}(t))~{}|~{}t=\-\infty]=0,$ (36)
for all $x\in\mathbb{R}^{n}$. By combining (35) and (36), $\mathrm{FRT}(\mathrm{Int}(S))\!=\!\{x~{}|~{}V_{\gamma}(x)>0\}$.

### v4

@@ 4.3 FRT Value Functions and CBFs
$=W(x)\quad(\text{when}\;V_{\gamma}(x)>0\;\text{or}\;W(x)>0)\vskip-15.00002pt$
Theorems 5 and 6 establish a tight connection between CBFs and FRTs:
Remark 2.
  1. 1.
    The robust control invariant set characterized by any CBF, $C_{W}$, is an FRT. In particular, this invariant set is the FRT of itself, i.e., $\mathrm{FRT}(C_{W})=C_{W}$ (by Corollary 1), and there may exist strict subsets $C$ whose FRT coincides with $C_{W}$.
  2. 2.
    For any CBF $W$ and any subset $C\subseteq C_{W}$, $W$ is a supersolution to the HJ-FRT-VI associated with $h_{C}$ satisfying $h_{C}(x)\leq W(x)$.
Theorem 6 and Remark 2 imply that any continuously differentiable supersolution of the HJ-FRT-VI is equivalent to a CBF. Therefore, general computational frameworks for supersolutions can serve as frameworks for generating CBFs, which we explore next. Through our formulation, we can design a robust control invariant set that outer-approximates the FRT, even when the initial set $C$ is not robustly control invariant.
5 Learning-based approach for FRT-CBF
We now propose a method that uses a neural network to learn a CBF whose control invariant set lies between the given constraint set $X$ and the initial set $C$. This framework is constructed by the mechanisms of Theorems 2 and 6 which verifies the learned control invariant set and the CBF as an outer- and over-approximation of the FRT and the FRT value function, respectively. We consider the CBF learned through this framework the neural forward reachable tube control barrier function (FRT-CBF) of the initial set $C$.
Assumption 3.
The given initial set $C$ is a subset of the viability kernel of the constraint set $X$.
The neural network we learn outputs two scalar functions, $u_{\theta}(x)$ and $v_{\theta}(x)$. Our objective is to learn the FRT-CBF, represented as
$W_{\theta}(x):=h_{C}(x)+u_{\theta}^{2}(x).$ (31)
By taking this representation, we can guarantee $W_{\theta}(x)\geq h_{C}(x)$ as a hard constraint.
The learning consists of three objectives:
  1. 1.
    Learn $W_{\theta}(x)$ that satisfies (30), so that according to Theorem (6), $C_{W}$ includes the minimal FRT.
  2. 2.
    Ensure that
    $W_{\theta}(x)\leq h_{X}(x)$
    is satisfied, so that the learned control invariant set $C_{W}$ is a subset of the constraint set $X$.
  3. 3.
    Ensure the control invaraince condition, i.e.,
    $\max_{u\in U}\min_{d\in D}\frac{\partial W_{\theta}}{\partial x}(x)\cdot f(x,u,d)\geq 0,\vskip-5.0pt$ (32)
    for all $x\in\partial C_{W}$, to satisfy the condition in Lemma 3.
To ensure the second condition, we use the second output of the neural network, $v_{\theta}(x)$, and impose a condition that
$W_{\theta}(x)=h_{C}(x)+u_{\theta}^{2}(x)=h_{X}(x)-v_{\theta}^{2}(x)\leq h_{X}(x).$ (33)
In summary, because $u_{\theta}(x)^{2}\geq 0$ and $v_{\theta}(x)^{2}\geq 0$, this parameterization enforces the relation
$h_{C}(x)\leq W_{\theta}(x)\leq h_{X}(x),$ (34)
and thus
$C\subset C_{W}\subset X.$
5.1 Related Work
Several recent works are closely related to the proposed training formulation. A number of studies apply the BRT-based CBVF formulation (in Section 2.3) to learn CBFs using neural networks [32, 47, 26]. These approaches are largely inspired by physics-informed machine learning for approximating solutions of PDEs, particularly HJ-PDEs arising in reachability analysis, as in [7]. Another line of work draws inspiration from backward reachability and learn CBFs that enlarge the control invariant set through Bellman backup-type updates [42].
These bodies of work may suffer from issues such as non-uniqueness, unboundedness, or discontinuity if the discount factor is not treated carefully. Although maximal invariant sets are an appealing property of backward reachability-based frameworks, our theory suggests that such formulations exhibit a fundamental discrepancy with well-behaved CBFs (e.g., Condition 2). This discrepancy can be addressed by instead adopting a forward reachability perspective.
Some works additionally incorporate pre-verified safe regions, similar to the initial set used in our formulation [17, 46]. These predefined safety regions typically serve as an initial guess for the CBF or as a constraint that prevents the learned invariant sets from becoming overly conservative. However, these works do not explicitly interpret the underlying problem through the lens of forward reachability.
Our work is related to these approaches in that we also use a neural network to approximate a supersolution of the HJ-FRT-VI, and the resulting loss function designs share similarities with those used in several neural CBF learning methods. However, crucially, we explicitly formulate the objective as learning a CBF whose zero-superlevel set outer-approximates the forward reachable tube (FRT) of a given initial set.
5.2 Loss Function
At each state sample $x_{i}$, we consider the following three loss terms for the learning.
1. HJ-FRT-VI Supersolution Loss: We evaluate the violation of the condition (30) with
$\ell_{\mathrm{frt}}(x_{i})=\min\Bigl\{\max_{u\in U}\min_{d\in D}\frac{\partial W_{\theta}}{\partial x}(x_{i})\cdot f(x_{i},u,d)+\gamma W_{\theta}(x_{i}),\;0\Bigr\}^{2}.$
This penalizes the violation of the inequality (29) that $W_{\theta}$ must satisfy. Note that (28) is already guaranteed by the representation (31). This term is active only when $W(x_{i})\geq 0$.
2. Set Inclusion Condition Loss: The condition (33) becomes equivalent to learning $u_{\theta}$ and $v_{\theta}$ such that
$u_{\theta}^{2}(x)+v_{\theta}^{2}(x)=h_{X}(x)-h_{C}(x).$ (35)
We evaluate the violation of the constraint (35) with
$\ell_{\mathrm{inclusion}}(x_{i})=\left((h_{X}(x_{i})-h_{C}(x_{i}))-\bigl(u_{\theta}(x_{i})^{2}+v_{\theta}(x_{i})^{2}\bigr)\right)^{2}$
so that the condition (34) is satisfied.
Figure 4: (a) Boundary of the learned FRT-CBF, which includes the initial set $C$ (dashed), and the $\mathrm{FRT}(\mathrm{Int}(C))$ (dash-dotted). The color of the boundary indicates the value of $\max_{u\in U}\min_{d\in D}\frac{\partial W_{\theta}}{\partial x}(x)\cdot f(x,u,d)$ which must be nonnegative for control invariance. The minimum value occured is -2.07e-4. (b) The learned $W_{\theta}(x)$ which is continuously differentiable, in comparison to $V_{\gamma}(x)$ and $h_{C}$, illustrating the relationship in Theorem 6.
3. Control Invariance Loss: We sample states near the boundary of $C_{W}$, finding a sample that satisfies
$0\leq W_{\theta}(x_{i})\leq\varepsilon,$ (36)
where $\varepsilon$ is set as a small threshold. We employ rejection sampling to sample the states that satisfy this condition. For these states, we evaluate the following loss
$\ell_{\mathrm{ci}}(x_{i})=\min\Bigl\{\max_{u\in U}\min_{d\in D}\frac{\partial W_{\theta}}{\partial x}(x_{i})\cdot f(x_{i},u,d),\;0\Bigr\}^{2}.$
This penalizes the state that does not satisfy the tangential condition in Lemma 3.
The final learning objective is
$\mathcal{L}=$ $\;w_{1}\mathbb{E}_{x\in\{x|W_{\theta}(x)\geq 0\}}[\ell_{\mathrm{frt}}(x_{i})]+w_{2}\mathbb{E}_{x}[\ell_{\mathrm{inclusion}}(x_{i})]$
$+w_{3}\mathbb{E}_{x\in\{x|0\leq W_{\theta}(x)\leq\varepsilon\}}[\ell_{\mathrm{ci}}(x_{i})].$ (37)
5.3 Neural Network & Training Procedure
The model is a multilayer perceptron with input dimension $n$, which takes in normalized state inputs, and output dimension $2$, outputing $(u_{\theta}(x),v_{\theta}(x))$. We use the softplus activation,
$\sigma(z)=\frac{1}{\beta}\log\!\bigl(1+e^{\beta z}\bigr),$ (38)
with $\beta>0$. This results in $u_{\theta},v_{\theta}$ to be continuously differentiable functions with differentiable level sets, and the maximum curvature can be controlled by choosing small enough $\beta$ [43].
By the chain rule, we compute the gradient of $W_{\theta}$ as
$\frac{\partial W_{\theta}(x)}{\partial x}$ $=\frac{\partial h_{C}(x)}{\partial x}+2u_{\theta}(x)\frac{\partial u_{\theta}(x)}{\partial x},$ (39)
where $\frac{\partial u_{\theta}(x)}{\partial x}$ is evaluated based on auto-differentiation.
Remark 3.
Since we represent the CBF as (31), for $W_{\theta}$ to satisfy Condition 2, we need to use $h_{C}$ that also satisfies Condition 2, meaning that our framework is currently limited to initial sets with a differentiable boundary. Similarly, the constraint target function $h_{X}$ also has to be a continuously differentiable function in order to satisfy (33). In practice, one can smoothen the given non-smooth initial and constraint set target functions.
Training details. We use a multilayer perceptron with three hidden layers with 512 features, and $\beta=10$. Each batch for one training iteration is generated by drawing 40,000 i.i.d. samples uniformly from the state range $10,10]\times[-10,10]$, and 5,000 samples satisfying (36) to evaluate the control invariance loss. We use the loss weights $w_{1}\!=\!w_{2}\!=\!w_{3}\!=\!1$. We used the Adam optimizer with learning rate $\eta=10^{-5}$, and run training for 10,000 iterations. The training takes around 30 minutes on NVIDIA RTX 5090 GPU.
5.4 Result
We revisit the double integrator example in Section 3.3, and focus on Case (a) where the chosen initial set $C$ is not control invariant.
The target function of the initial set $h_{C}$ is a circular quadratic function centered at the origin with radius $r=4$:
$h_{C}(x)=1-\frac{\left\lVert x\right\rVert^{2}}{r^{2}}$
The constraint function $h_{X}$ is set as a circular quadratic function centered at the origin with radius $R=9$:
$h_{X}(x)=1-\frac{\left\lVert x\right\rVert^{2}}{R^{2}}.$
The training loss at the last iteration of the training is 2.25e-7, implying that the CBF is learned successfully with negligible errors. The learned control invariant set and the FRT-CBF is visualized in Figure 4. Together, we also visualize the FRT value function $V_{\gamma}(x)$, computed by the finite horizon-based method described in Appendix 8.8. We used $\gamma=2$ for both computations.
@@ 6 Discussion
Although inspired by existing works that establish the connection between reachability and robust control invariance, our work is the first paper that connects forward reachability to the analysis of robust control invariance. Existing works have focused on backward reachability and proposed various value functions that characterize the inevitable BRT of the failure region, for both finite time [13, 31, 33, 22] and infinite time [21, 49, 2, 44]. However, none of these functions are CBFs, as discussed in Section 2.3 and summarized in Table 1. The main issue with these formulations is that the value function is either flat inside the invariant set (when $\lambda>0$ in (14)), or can become discontinuous (when $\lambda\leq 0$), both of which make its use as a CBF infeasible.
Our formulation of forward reachability ensures compliance with the barrier constraint in the CBF definition. Additionally, the value function is both continuous and bounded in $\mathbb{R}^{n}$, while the corresponding HJ-PDE has a unique solution. The central idea behind our approach is the usage of the discount factor backward in time in the definition of the discounted FRT value function (18). In contrast to the discount in BRT formulations leading to the emergence of $\gamma V_{\gamma}(x)$ in the corresponding HJ-PDEs (Table 1, first row), the usage of discount in this way leads to the emergence of positive $\gamma V_{\gamma}(x)$ term in the HJ-FRT-VI (21), and thus the satisfaction of the barrier constraint (Table 1, last row). Moreover, $e^{\gamma t}$ vanishes as $t$ approaches -$\infty$, thereby ensuring continuous, bounded value functions and the solution uniqueness of the HJ-FRT-VI, resulting from the contraction mapping property outlined in Section 4.2.
@@ 7 Conclusions
One salient condition underlying our study is the differentiability of the boundaries of control invariant sets. Although this condition is required for the zero-superlevel sets of any valid CBFs, recent extensions of CBFs can alleviate this condition [16, 23]. A deeper understanding of the implications of this condition is crucial for broadening the applicability of our results. For example, the relevance of such regularity properties of the boundary of the safe set for stochastic systems is elucidated in [11]. Finally, application of the proposed neural FRT-CBF framework to more practical high-dimensional systems is left as an important future research direction.
@@ 8 Appendix
8.1 Proof of Theorem 1
For the proof of Theorem 1, we have to reason about the backward flow of the dynamics, as in (16). Thus, we first consider the notion of backward control invariance, the mirrored version of the forward control invariance.
Definition 7 (Robustly backward control invariant).
Corollary 3.
A set $S$ satisfying Condition 1 is robustly backward control invariant if and only if for all $x\in\partial S$,
$\exists u\in U~\textnormal{s.t.}~-\frac{\partial h_{S}}{\partial x}(x)\cdot f(x,u,d)\geq 0\;\forall d\in D,\vskip-10.00002pt$ (40)
for $h_{S}$ satisfying Condition 2.
Lemma 5.
A set $S$ satisfying Condition 1 is robustly forward control invariant if and only if $\mathrm{Int}(S)^{c}$ is robustly backward control invariant.
Proof.
By Lemma 3, $S$ is robustly forward control invariant if and only if for all $x\in\partial S$, (10) is satisfied. Note that $\partial S=\partial\mathrm{Int}(S)^{c}$ and $\mathrm{Int}(S)^{c}$ and $h_{\mathrm{Int}(S)^{c}}:=-h_{S}$ also satisfy Conditions 1 and 2, respectively. Since $\frac{\partial h_{\mathrm{Int}(S)^{c}}}{\partial x}(x)=-\frac{\partial h_{S}}{\partial x}(x)$, (10) is equivalent to
In Lemma 5, Condition 1 guarantees that, for a state $x_{1}$ on the boundary of $S$, if there exists a particular control $u_{1}$ such that $f(x_{1},u_{1},d)$ points inward to $S$ for all $d\in D$, $f(x_{1},u_{1},d)$ points outwards to $S$ for all $d\in D$.
Definition 8.
[…]
This definition applies the concept of leadership kernel in [12] to the backward dynamics. Given any closed set $C\subset\mathbb{R}^{n}$, the $\mathrm{VK}^{-}(C)$ is the largest robustly backward invariant set in $C$.
Lemma 6.
$\{\mathrm{FRT}(C)\}^{c}=\mathrm{VK}^{-}(C^{c}).\vskip-5.0pt$ (42)
Proof.
The complement relationship (42) follows directly from Definitions 6 and 8. That $\mathrm{VK}^{-}(C)$ is a closed set is proven in [12], and $\mathrm{FRT}(C)$ being an open set results from the complement relationship. ∎
Based on the lemmas above, we present the proof of Theorem 1 below.
Proof.
(Theorem 1) Since $\mathrm{VK}^{-}(C^{c})$ is the minimal robustly backward invariant set contained in $C^{c}$, Lemmas 5 and 6 imply that its complement $\mathrm{FRT}(C)$ is the maximal robustly forward control invariant set containing $C$.
8.2 Proof of Corollary 1
By Lemma 5, a set $S$ satisfying Condition 1 is robustly control invariant if and only if the complement of $\mathrm{Int}(S)$ is robustly backward control invariant. By Definition 8, this is equivalent to $\mathrm{VK}^{-}(\{\mathrm{Int}(S)\}^{c})=\{\mathrm{Int}(S)\}^{c}$. By Lemma 6, this condition is further equivalent to $\mathrm{FRT}(\mathrm{Int}(S))=\mathrm{Int}(S)$.
8.3 Proof of Theorem 2
We want to prove that $\mathrm{FRT}(\mathrm{Int}(C))\subset S$, which is sufficient to prove $\{\mathrm{Int}(S)\}^{c}\subset\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}$. Since the set $S$ satisfies Condition 1 and is robustly control invariant, by Lemma 5, $\{\mathrm{Int}(S)\}^{c}$ is robustly backward control invariant. Also, since $C\subset S$, $\{\mathrm{Int}(S)\}^{c}\subset\{\mathrm{Int}(C)\}^{c}$. By Lemma 6 $\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}=\mathrm{VK}^{-}(\{\mathrm{Int}(C)\}^{c})$, which is the largest robustly backward invariant in $\{\mathrm{Int}(C)\}^{c}$. Therefore, $\{\mathrm{Int}(S)\}^{c}\subset\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}$.
@@ 8.4 Proof of Lemma 4
$\rho_{\mathrm{u}}(\mathrm{u})(-t)=\mathrm{u}(t),\quad\forall t\in[0,\infty);\vskip-5.0pt$ (43)
@@
$\rho_{\xi_{d}}(\xi_{d})[\mathrm{u}^{-}](-t)=\xi_{d}[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t),\vskip-5.0pt$ (44)
@@
$\mathrm{x}^{-}(-t)=\mathrm{x}(t)\quad\forall t\in[0,\infty),\vskip-5.0pt$ (45)
@@
$\dot{\mathrm{x}}(t)=-f(\mathrm{x}(t),\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})(t),\rho_{\xi_{d}}^{-1}(\Xi_{d}^{-})[\rho_{\mathrm{u}}^{-1}(\mathrm{u}^{-})](t))\vskip-2.5pt$
for $t>0$, and $\mathrm{x}(0)=x$. Since $\rho_{\mathrm{u}}$ and $\rho_{\xi_{d}}$ are one-to-one, the viability kernel of $\mathrm{Int}(C)^{c}$ under f-f, defined in Definition 8, is equivalent to the following set:
$\mathrm{VK}^{-}(C)=$ $\{x\in\mathbb{R}^{n}~|~\forall\xi_{d}\in\Xi_{d},\varepsilon>0,T>0,\exists\mathrm{u}\in\mathcal{U}\text{ s.t.}$
$\forall t\in[0,T],\mathrm{x}(t)\in C+B_{\varepsilon}\},\vskip-5.0pt$ (46)
@@
$\dot{\mathrm{x}}(t)=-f(\mathrm{x}(t),\mathrm{u}(t),\xi_{d}[\mathrm{u}](t)),\;\forall t>0,\;\;\mathrm{x}(0)=x.$
By [49, Lemma 1], the viability kernel (46) of $\mathrm{Int}(C)^{c}$ is characterized by a particular value function:
$\mathrm{VK}^{-}(\mathrm{Int}(C)^{c})\!=\!\Big\{x~\Big|~\!\inf_{\xi_{d}\in\Xi_{d}}\sup_{\mathrm{u}\in\mathcal{U}}\inf_{t\in[0,\infty)}e^{-\gamma t}\big(\!-\!h_{C}(\mathrm{x}(t))\big)\!=\!0\Big\},$
where $\mathrm{x}$ solves (47). By (45) and one-to-one properties of $\rho_{\mathrm{u}}$ and $\rho_{\xi_{d}}$$\mathrm{VK}^{-}(\mathrm{Int}(C)^{c})=\{x~|~V_{\gamma}(x)=0\}$. By Lemma 6,
$\mathrm{FRT}(\mathrm{Int}(C))=\{x~|~V_{\gamma}(x)\neq 0\}.\vskip-5.0pt$ (48)
Since $h_{C}$ is bounded,
$V_{\gamma}(x)\geq\sup_{\xi_{d}^{-}\in\Xi_{d}^{-}}\inf_{\mathrm{u}^{-}\in\mathcal{U}^{-}}[e^{\gamma t}h_{C}(\mathrm{x}^{-}(t))~|~t=\-\infty]=0,$ (49)
for all $x\in\mathbb{R}^{n}$. By combining (48) and (49), $\mathrm{FRT}(\mathrm{Int}(C))\!=\!\{x~|~V_{\gamma}(x)>0\}$.
This also implies that $\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}=\{x\mid V_{\gamma}(x)\leq 0\}$. From the definition of the FRT value function in (18), we have $V_{\gamma}(x)\geq h_{C}(x)$ for all $x$, since $J_{\gamma}(x,\mathrm{u}^{-},\xi_{d}^{-})\geq h_{C}(e^{\gamma 0}\mathrm{x}^{-}(0))=h_{S}(x)$. Therefore, it follows that $\{\mathrm{FRT}(\mathrm{Int}(C))\}^{c}=\{x\mid V_{\gamma}(x)=0\}$.